OpenStudy (anonymous):

if x + 1/x = 5 how about x^3 + 1/ x^3 ?

OpenStudy (saifoo.khan):

o.O

OpenStudy (anonymous):

the question is this \[x^{3} + 1/ x^{3}\]

OpenStudy (anonymous):

cube x + 1/x

OpenStudy (anonymous):

(x + 1/x )^3 = 125 = x^3 + 1/x^3+ 3x + 3/x 125 = x^3 + 1/x^3 + 3(x+1/x) 125 = x^3 + 1/x^3 +15 hence x^3 + 1/x^3 = 110

OpenStudy (anonymous):

cube ?

OpenStudy (anonymous):

Thanks !